### v2

@@ Abstract
Three-dimensional (3D) spot beamfocusing (SBF), in contrast to conventional angular-domain beamforming, concentrates radiating power within a very small volume in both radial and angular domains in the near-field zone. Recently the implementation of channel-state-information (CSI)-independent machine learning (ML)-based approaches have been developed for effective SBF using extremely-largescale-programable-metasurface (ELPMs). These methods involve dividing the ELPMs into subarrays and independently training them with Deep Reinforcement Learning to jointly focus the beam at the desired focal point (DFP). This paper explores near-field SBF using ELPMs, addressing challenges associated with lengthy training times resulting from independent training of subarrays. To achieve a faster CSI-independent solution, inspired by the correlation between the beamfocusing matrices of the subarrays, we leverage transfer learning techniques. First, we introduce a novel similarity criterion based on the phase distribution image (PDI) of subarray apertures. Then we devise a subarray policy propagation scheme that transfers the knowledge from trained to untrained subarrays. We further enhance learning by introducing quasi-liquid layers as a revised version of the adaptive policy reuse technique. We show through simulations that the proposed scheme improves the training speed about 5 times. Furthermore, for dynamic DFP management, we devised a DFP policy blending process, which augments the convergence rate up to 8-fold.
@@ I Introduction
Electromagnetic radiating apertures, encompassing both continuous (e.g., parabolic reflectors) and discrete (e.g., planar arrays), exhibit three distinct radiation regions categorized by their distance from the aperture, as detailed in [1]. The region closest to the aperture is the reactive near-field, characterized by the dominance of reactive power over radiating power. The furthest region, known as the far-field or Fraunhofer zone, exhibits planar propagation of radiating waves, resulting in a distance-independent radiation pattern, as depicted in (Fig. 1-a) In this zone, the radiating pattern of the aperture, relative to an isotropic antenna, can only vary within the angular domain. The Fresnel zone, situated between the reactive near-field and the far-field zones, is characterized by a spherical wavefront. This inherent property enables beamfocusing (as illustrated in (Fig. 1-b), where manipulation of the phase at distinct aperture points allows concentration of radiating power into a small focal region [2]. The Fresnel zone for a continuous aperture is characterized by the region where the distance from the aperture (denoted by $r$) holds in the range of $\underline{D}\leq r\leq$ $\overline{D}$ where $D$ is the largest dimension of the aperture, $\lambda$ is the wavelength, $\overline{D}=2D^{2}/\lambda$ for the on-boresight scenario and $\underline{D}$ is limited to the space very close to the antenna aperture [3, 4].
@@
Spot beamfocusing is defined by the precise control of the focal region to an exceptionally small size, thereby concentrating the bulk of radiating power into a localized, spot-like area [23, 24] (Fig. 1-c). To achieve SBF, a high ratio of $D/\lambda$ is required [2], which can be accomplished through decreasing $\lambda$ as well as increasing $D$. The former is achieved through transitioning to mm-wave/THz frequencies. The latter can be realized by employing extremely large-scale array antennas (ELAAs). However, implementing SBF with CPAs necessitates bulky and expensive ELAAs. Additionally, holographic MIMO surfaces have limitations in fully adjustable DFP and power transmission [19]. A viable option for SBF implementation is to use programmable metasurfaces, which provide low-cost, mass-producible, and reconfigurable ELAAs. Realizing SBF with ELAAs necessitates coherent signal arrival from all elements at the DFP. his necessitates phase shifters for each element, demanding the precise computation of an optimal phase-shift matrix (which is called the beamfocusing matrix in the paper). While various channel-state-information (CSI)-independent far-field beam training methods have been proposed for multi-subarray massive MIMO systems, such as the low-complexity scheme exploiting the spatial correlation between ELAA elements in [25] and the fast multibeam training method for sub-connected hybrid beamforming [26], the application of these methods to near-field SBF are practically restrictive and hence other techniques should be developed. The challenges stem from estimating the CSI related to a massive number of array elements and the intricate nature of near-field channel matrices (characterized by spherical wave propagation and non-sparse angular-domain representation) [27, 28]. Consequently, recent research has explored machine learning (ML)-based CSI-independent SBF structures, as exemplified in [23].
@@ I-A Motivation and Contributions
  • Novel Correlation-Aware TL Approach: A key contribution of this paper is the introduction of a novel similarity metric based on the phase distribution image (PDI) of subarray apertures, wherein the excitation phases of array elements are considered as the pixels of a 2D image. The proposed metric builds a correlation assessment between different PDIs, facilitating efficient knowledge transfer from the trained agents to untrained ones. Building upon this correlation assessment, we propose a novel TL-based subarray policy propagation scheme, which increases the speed of finding the optimal beamfocusing matrix 4-fold compared to the case where no TL is leveraged, as evidenced through numerical results.
  • Quasi-Liquid Layers (QLL): To improve training efficiency, we introduce a revised adaptive policy reuse technique called quasi-liquid layer (QLL). This involves adding extra layers to deep neural networks (DNNs) for actor and critic networks. The inner and middle layers are kept frozen, while the outer layers gradually become liquid. More exactly, the learning rates for the outer layers are adjusted from zero (frozen) to one (liquid). This procedure results in an extra 20% increase in training speed compared to the subarray policy propagation scheme without QLL.
@@ II-A System Model
$x=\boldsymbol{1}^{\mathrm{T}}_{N_{r}}\left(\boldsymbol{w}^{*}\odot\boldsymbol{h}\right)\boldsymbol{1}_{N_{c}}+\nu$ (3)
@@ V-A1 Subarray Policy Propagation
At the initialization phase, we assume that all subarrays of the ELPM constitute the set of students $\mathcal{M}^{s}\leftarrow\mathcal{M}$, and there exists no pre-trained teacher subarray, i.e., $\mathcal{M}^{t}\leftarrow\O$. To be able to transfer the policies, we need at least $K_{1}\geq 1$ trained subarrays’ policies ($|\mathcal{M}^{t}|\geq K_{1}$). In our simulations, we have selected $K_{1}=1$. This choice accounts for the maximum potential benefit of any possible teacher candidate at the initialization phase, even when the set of teacher candidates is of a small cardinality. Following the initialization phase, a teacher policy library is established. This library’s knowledge is dynamically applied through policy reuse techniques to efficiently train student policies. The library grows dynamically and becomes enriched over time. This process is achieved in the Main Procedure in lines 3-21 of the algorithm 1. The policy reuse technique requires finding the best source policy for the target task. We apply the similarity criterion on PDIs to find the best teacher policy for training a specific student subarray policy. In doing so, in steps 4-5, first, a student candidate $m^{s}$ is selected from $\mathcal{M}^{s}$ as some neighbor of one of the teacher candidates $\mathcal{M}^{t}$, and then the candidate teachers $\mathcal{M}^{t,cnd}$ is selected as the first $K_{2}$ trained agents of subarrays with geometrical distance from the student subarray location in the ELPM. To find the best teacher candidate for $m^{s}$, for each $m\in\mathcal{M}^{t,cnd}$, first we start a test mode by temporarily transferring the policy $\pi^{(m)}$ to $\pi^{(m^{s})}$ and then begin a temporary training phase with a very limited number of iterations for subarray $m^{s}$ leading to the temporary PDI $\boldsymbol{w}^{(m^{s,t})}$. Here, we consider that the temporary phase learning rate $l^{r,t}$, and the exploration noise decay rate $\alpha^{xp,t}$ are much higher respectively than those relating to the final training phase ($l^{r}$ and $\alpha^{xp}$). This ensures that the estimated results are obtained from the test mode in very few numbers of iterations. From all teacher candidates, we select the one corresponding to the highest ECC according to step 14 of algorithm 1. The search for the optimal teacher candidate can be terminated if the ECC value exceeds a certain threshold. In that case, the current candidate is regarded as the best teacher. This is seen in steps 10-12 of the algorithm 1. After selecting the best teacher, we use the teacher policy for training the student subarray. Direct policy reuse, where the identical policy from a teacher policy is mechanically applied to a student policy without adjustments, is demonstrably suboptimal. This arises due to the inherent non-equivalence of the source and target tasks’ optimal beamfocusing submatrices. This disparity stems from the varying physical locations of the subarrays within the ELPM, necessitating tailored policies for optimal performance. Therefore, we use a revised adaptive policy reuse technique called QLLs, which is described in Section V-A2. At this stage, we employ the teacher policy $\pi^{(m^{t})}$ for training the student policy $\pi^{(m^{s})}$ according to step 16 of the algorithm 1. Then, the trained student subarray $m_{s}$ is removed from $\mathcal{M}^{s}$ and added to the set of possible teacher candidates $\mathcal{M}^{t}$. This enriches the policy library and increases the flexibility of finding more appropriate teachers for training the rest of the students. Since this learning process gradually propagates throughout the ELPM subarrays, we have entitled it as subarray policy propagation. The solution of P1 is finally obtained by constituting the beamfocusing matrix $\widehat{\boldsymbol{w}}$ in (2) using $\widehat{\boldsymbol{w}}^{(m)},\forall m\in\mathcal{M}$. The proposed subarray policy propagation algorithm accomplishes the training process for all $M$ subarrays of ELPM, with a much higher convergence speed than that achieved in [23]. More specifically, we will show in the numerical results that using the subarray policy propagation technique enhances the average training speed of subarrays by about 4 times.
@@ V-B2 How to Select Policy Components
Depending on accessible technologies and available parameters, different strategies might be adopted to select the set of component policies $\boldsymbol{\pi}_{\mathrm{cmp}}$ (corresponding to the focal points $\boldsymbol{\mathcal{D}}_{\mathrm{cmp}}$) from the policy library. For example, if the location of the focal points in $\boldsymbol{\mathcal{D}}^{+}$ is available through localization schemes, we can take up to $K$ focal points with the closest distance to $\mathcal{D}^{-}$, as long as each distance is less than a minimum threshold value; this raises the probability of higher similarity of the PDIs of $\boldsymbol{\mathcal{D}}_{\mathrm{cmp}}$ with that of $\mathcal{D}^{-}$. For the case where the exact estimation of locations is not available, we may simply choose the last $K$ policies recently added to the library. This scheme is valid when the duration for the DFP to transition from one position to the subsequent one is not less than the time required to train the subarray policies at each respective DFP.

### v3

@@ Abstract
Three-dimensional (3D) spot beamfocusing (SBF), in contrast to conventional angular-domain beamforming, concentrates radiating power within a very small volume in both radial and angular domains in the near-field zone. Recently the implementation of channel-state-information (CSI)-independent machine learning (ML)-based approaches have been developed for effective SBF using extremely large-scale programmable metasurface (ELPMs). These methods involve dividing the ELPMs into subarrays and independently training them with Deep Reinforcement Learning to jointly focus the beam at the desired focal point (DFP). This paper explores near-field SBF using ELPMs, addressing challenges associated with lengthy training times resulting from independent training of subarrays. To achieve a faster CSI-independent solution, inspired by the correlation between the beamfocusing matrices of the subarrays, we leverage transfer learning techniques. First, we introduce a novel similarity criterion based on the phase distribution image (PDI) of subarray apertures. Then we devise a subarray policy propagation scheme that transfers the knowledge from trained to untrained subarrays. We further enhance learning by introducing quasi-liquid layers as a revised version of the adaptive policy reuse technique. We show through simulations that the proposed scheme improves the training speed about 5 times. Furthermore, for dynamic DFP management, we devised a DFP policy blending process, which augments the convergence rate up to 8-fold.
@@ I Introduction
Electromagnetic radiating apertures, encompassing both continuous (e.g., parabolic reflectors) and discrete (e.g., planar arrays), exhibit three distinct radiation regions categorized by their distance from the aperture, as detailed in [1]. The region closest to the aperture is the reactive near-field, characterized by the dominance of reactive power over radiating power. The furthest region, known as the far-field or Fraunhofer zone, exhibits planar propagation of radiating waves, resulting in a distance-independent radiation pattern, as depicted in (Fig. 1-a) In this zone, the radiating pattern of the aperture, relative to an isotropic antenna, can only vary within the angular domain. The Fresnel zone, situated between the reactive near-field and the far-field zones, is characterized by a spherical wavefront. This inherent property enables beamfocusing (as illustrated in (Fig. 1-b), where manipulation of the phase at distinct aperture points allows concentration of radiating power into a small focal region [2]. The Fresnel zone for a continuous aperture is characterized by the region where the distance from the aperture (denoted by $r$) holds in the range of $\underline{D}\leq r\leq$ $\overline{D}$ where $D$ is the largest dimension of the aperture, $\lambda$ is the wavelength, $\overline{D}=2D^{2}/\lambda$ for the on-boresight scenario and $\underline{D}$ is limited to the space very close to the antenna aperture [3]. The authors of [4, 5] have shown that for the off-boresight observation point, $\overline{D}$ can be extended to about 4 times that relating to the on-boresight case.
@@
Spot beamfocusing is defined by the precise control of the focal region to an exceptionally small size, thereby concentrating the bulk of radiating power into a localized, spot-like area [24, 25] (Fig. 1-c). To achieve SBF, a high ratio of $D/\lambda$ is required [2], which can be accomplished through decreasing $\lambda$ as well as increasing $D$. The former is achieved through transitioning to mm-wave/THz frequencies. The latter can be realized by employing extremely large-scale array antennas (ELAAs). However, implementing SBF with CPAs necessitates bulky and expensive ELAAs. Additionally, holographic MIMO surfaces have limitations in fully adjustable DFP and power transmission [20]. A viable option for SBF implementation is to use programmable metasurfaces, which provide low-cost, mass-producible, and reconfigurable ELAAs. Realizing SBF with ELAAs necessitates coherent signal arrival from all elements at the DFP. This necessitates phase shifters for each element, demanding the precise computation of an optimal phase-shift matrix (which is called the beamfocusing matrix in the paper). While various channel-state-information (CSI)-independent far-field beam training methods have been proposed for multi-subarray massive MIMO systems, such as the low-complexity scheme exploiting the spatial correlation between ELAA elements in [26] and the fast multibeam training method for sub-connected hybrid beamforming [27], the application of these methods to near-field SBF are practically restrictive and hence other techniques should be developed. The challenges stem from estimating the CSI related to a massive number of array elements and the intricate nature of near-field channel matrices (characterized by spherical wave propagation and non-sparse angular-domain representation) [28, 29]. Consequently, recent research has explored machine learning (ML)-based CSI-independent SBF structures, as exemplified in [24].
@@ I-A Motivation and Contributions
  • Novel Correlation-Aware TL Approach: A key contribution of this paper is the introduction of a novel similarity metric, based on the phase distribution image (PDI) of subarray apertures, wherein the excitation phases of array elements are considered as the pixels of a 2D image. The proposed metric builds a correlation assessment between different PDIs, facilitating efficient knowledge transfer from the trained agents to untrained ones. Building upon this correlation assessment, we propose a novel TL-based subarray policy propagation scheme, which increases the speed of finding the optimal beamfocusing matrix 4-fold compared to the case where no TL is leveraged, as evidenced through numerical results.
  • Quasi-Liquid Layers: To improve training efficiency, we introduce a revised adaptive policy reuse technique called quasi-liquid layer (QLL). This involves adding extra layers to deep neural networks (DNNs) for actor and critic networks. The inner and middle layers are kept frozen, while the outer layers gradually become liquid. More exactly, the learning rates for the outer layers are adjusted from zero (frozen) to one (liquid). This procedure results in an extra 20% increase in training speed compared to the subarray policy propagation scheme without QLL.
@@ II-A System Model
$x=\boldsymbol{1}^{\mathrm{T}}_{N_{r}}\left(\boldsymbol{w}^{*}\odot\boldsymbol{h}\right)\boldsymbol{1}_{N_{c}}s+\nu$ (3)
@@ V-A1 Subarray Policy Propagation
At the initialization phase, we assume that all subarrays of the ELPM constitute the set of students $\mathcal{M}^{s}\leftarrow\mathcal{M}$, and there exists no pre-trained teacher subarray, i.e., $\mathcal{M}^{t}\leftarrow\O$. To be able to transfer the policies, we need at least $K_{1}\geq 1$ trained subarrays’ policies ($|\mathcal{M}^{t}|\geq K_{1}$). In our simulations, we have selected $K_{1}=1$. This choice accounts for the maximum potential benefit of any possible teacher candidate at the initialization phase, even when the set of teacher candidates is of a small cardinality. Following the initialization phase, a teacher policy library is established. This library’s knowledge is dynamically applied through policy reuse techniques to efficiently train student policies. The library grows dynamically and becomes enriched over time. This process is achieved in the Main Procedure in lines 3-21 of the algorithm 1. The policy reuse technique requires finding the best source policy for the target task. We apply the similarity criterion on PDIs to find the best teacher policy for training a specific student subarray policy. In doing so, in steps 4-5, first, a student candidate $m^{s}$ is selected from $\mathcal{M}^{s}$ as some neighbor of one of the teacher candidates $\mathcal{M}^{t}$, and then the candidate teachers $\mathcal{M}^{t,cnd}$ is selected as the first $K_{2}$ trained agents of subarrays with geometrical distance from the student subarray location in the ELPM. To find the best teacher candidate for $m^{s}$, for each $m\in\mathcal{M}^{t,cnd}$, first we start a test mode by temporarily transferring the policy $\pi^{(m)}$ to $\pi^{(m^{s})}$ and then begin a temporary training phase with a very limited number of iterations for subarray $m^{s}$ leading to the temporary PDI, $\boldsymbol{w}^{(m^{s,t})}$. Here, we consider that the temporary phase learning rate $l^{r,t}$, and the exploration noise decay rate $\alpha^{xp,t}$ are much higher respectively than those relating to the final training phase ($l^{r}$ and $\alpha^{xp}$). This ensures that the estimated results are obtained from the test mode in very few numbers of iterations. From all teacher candidates, we select the one corresponding to the highest ECC according to step 14 of algorithm 1. The search for the optimal teacher candidate can be terminated if the ECC value exceeds a certain threshold. In that case, the current candidate is regarded as the best teacher. This is seen in steps 10-12 of the algorithm 1. After selecting the best teacher, we use the teacher policy for training the student subarray. Direct policy reuse, where the identical policy from a teacher policy is mechanically applied to a student policy without adjustments, is demonstrably suboptimal. This arises due to the inherent non-equivalence of the source and target tasks’ optimal beamfocusing submatrices. This disparity stems from the varying physical locations of the subarrays within the ELPM, necessitating tailored policies for optimal performance. Therefore, we use a revised adaptive policy reuse technique called QLLs, which is described in Section V-A2. At this stage, we employ the teacher policy $\pi^{(m^{t})}$ for training the student policy $\pi^{(m^{s})}$ according to step 16 of the algorithm 1. Then, the trained student subarray $m_{s}$ is removed from $\mathcal{M}^{s}$ and added to the set of possible teacher candidates $\mathcal{M}^{t}$. This enriches the policy library and increases the flexibility of finding more appropriate teachers for training the rest of the students. Since this learning process gradually propagates throughout the ELPM subarrays, we have entitled it as subarray policy propagation. The solution of P1 is finally obtained by constituting the beamfocusing matrix $\widehat{\boldsymbol{w}}$ in (2) using $\widehat{\boldsymbol{w}}^{(m)},\forall m\in\mathcal{M}$. The proposed subarray policy propagation algorithm accomplishes the training process for all $M$ subarrays of ELPM, with a much higher convergence speed than that achieved in [24]. More specifically, we will show in the numerical results that using the subarray policy propagation technique enhances the average training speed of subarrays by about 4 times.
@@ V-B2 How to Select Policy Components
Depending on accessible technologies and available parameters, different strategies might be adopted to select the set of component policies $\boldsymbol{\pi}_{\mathrm{cmp}}$ (corresponding to the focal points $\boldsymbol{\mathcal{D}}_{\mathrm{cmp}}$) from the policy library. For example, if the location of the focal points in $\boldsymbol{\mathcal{D}}^{+}$ is available through localization schemes, we can take up to $K$ focal points with the closest distance to $\mathcal{D}^{-}$, as long as each distance is less than a minimum threshold value; this raises the probability of higher similarity of the PDIs of $\boldsymbol{\mathcal{D}}_{\mathrm{cmp}}$ with that of $\mathcal{D}^{-}$. For the case where the exact estimation of locations is not available, we may simply choose the last $K$ policies recently added to the library. This scheme is applicable when the duration for the DFP to transition from one position to the subsequent one is not less than the time required to train the subarray policies at each respective DFP.